What is the model for unkowns for this problem: Mrs. B invested $30,000: part at 5%, and part at 8%. The total interest on the investment was $2100. How much did she invest at each rate?

I started here: $30,000-x=7
x=.05
y=.08
Stuck now...

Answer by scott8148(6628)

x + y = 30000 ___ x = 30000 - y

(5%)x + (8%)y = 2100

substituting ___ (.05)(30000 - y) + (.08y) = 2100 ___ 1500 + .03y = 2100

solve for y; then substitute back to find x